how do i find the equation of a line through two points, (4,3) and (2,4)? Please explain, the book says the answer is y=-1/2x+5 but I don't get how they got -1/2 and 5.
Found 2 solutions by Greenfinch, richwmiller:
Answer by Greenfinch(383)

First, the gradient.
The x co-ord goes from 4 to 2 which is -2
The y co-ord goes from 3 to 4 which is +1
gradient is y/x = 1/-2 = -1/2.
The value of y when x = 0 will be 5 the point (0,5) known as the y intercept
The equstion is y = gradient x + y intercept
which is y = - 1/2x + 5

Answer by richwmiller(17219)

-1/2=m
5=b
y=mx+b
y-y1=m*(x-x1)
where m=slope the -1/2
(4-3)=m*(2-4)
1/-2=-1/2=m
y=mx+b
now choose either point.
4=-1/2*2+b
4=-1+b
5=b